OpenStudy (anonymous):

How many solutions are there for the equation? x2 + 16 = 8x

OpenStudy (anonymous):

2 solution because its X^2

OpenStudy (callisto):

x^2 + 16 = 8x x^2 -8x+ 16 = 0 delta = (-8)^2 - 4(1)(16) = 64 - 64 =0 Since delta = 0, there is a double real root.

OpenStudy (anonymous):

okay, thanks guys :)